OpenStudy (anonymous):

PLZZZZZ HELPP!!!! Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n^2 - n + 2

OpenStudy (helder_edwin):

\[ \large 2\mid(n^2-n+2) \]

OpenStudy (helder_edwin):

if n=1 then \[ \large n^2-n+2=1^2-1+2=2 \] which is a multiple of 2.

OpenStudy (helder_edwin):

let's assume that n>1 and \[ \large 2\mid(n^2-n+2) \] we have to prove that \[ \large 2\mid[(n+1)^2-(n+1)+2] \].

OpenStudy (helder_edwin):

Let's see \[ \large (n+1)^2-(n+1)+2=n^2+2n+1-n-1+2 \] \[ \large =(n^2-n+2)+2n=2\cdot\alpha+2n=2(\alpha+n) \]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

where did the a come from?

OpenStudy (helder_edwin):

don't forget that \[ \large 2\mid(n^2-n+2)\quad\Leftrightarrow\quad n^2-n+2=2\cdot\alpha \] this the induction hipothesis.

OpenStudy (anonymous):

oh yes i forgot.

OpenStudy (anonymous):

so that is it??

OpenStudy (helder_edwin):

yes!!! do u know how to do induction??

OpenStudy (anonymous):

kind of. but you really helped!! thank u soo much!!

OpenStudy (helder_edwin):

u r welcome glad to help